# Parameter-Efficient Fine-Tuning via Circular Convolution

### v1

@@ 3.5 Complexity Analysis
We compare time compelxity and space complexity of LoRA, VeRA and C3A in Table 1. Detailed analysis follows.
Table 1: Time and space complexity comparison of LoRA, VeRA and C3A. We split the space complexity into Parameter number and Other auxiliary tensors to help better understand the differences. We highlight that in practice, to achieve similar performance, $b\leq r\ll r_{v}$.
Method Time # Parameter # Other # Total
LoRA $\mathcal{O}(rd)$ $2rd$ $r$ $2rd+r$
VeRA $\mathcal{O}(r_{v}d)$ $r_{v}+d$ $2r_{v}d+r_{v}+d$ $2(r_{v}d+r_{v}+d)$
C3A $\mathcal{O}(bd\log\frac{d}{b})$ $bd$ 00 $bd$
@@ 3.5.1 Time Complexity
LoRA incorporates low-rank matrices $\mathbf{A}$ and $\mathbf{B}$, which are sequentially multiplied with the activation vector, resulting in a time complexity of $\mathcal{O}(rd)$. Typically, $r\ll d$. Due to the $\mathcal{O}(d\log d)$ complexity of the FFT algorithm used in Equation 1, C3A achieves a time complexity of $\mathcal{O}(bd\log\frac{d}{b})$. Given that $b\ll d$, C3A is comparable to LoRA in terms of time complexity. Conversely, VeRA, despite its high-rank nature and minimal number of trainable parameters, incurs a prohibitive time complexity of $\mathcal{O}(r_{v}d)$, where $r_{v}$ can exceed $d$. Therefore, striking a balance between high rank and computational efficiency is never a low-hanging fruit.
@@ 3.5.2 Space Complexity
We analyze the space complexity of LoRA, VeRA, and C3A during training. The differences among these methods primarily arise from the trainable parameter size and the auxiliary tensors required for the forward pass and backpropagation. For LoRA, an intermediate tensor of size $r$ is introduced (i.e., $\mathbf{Ax}\in\mathbb{R}^{r}$). In contrast, VeRA necessitates two random projection matrices, with a total size of $2r_{v}d$. Additionally, VeRA introduces two intermediate tensors of sizes $d$ and $r_{v}$. Thanks to the efficient backpropagation algorithm detailed in Section 3.3, C3A does not require any additional intermediate tensors. This optimization significantly reduces its space complexity compared to LoRA and VeRA.
@@ 4 Experiment
In this section, we comprehensively evaluate C3A in both natural language processing (NLP) and computer vision (CV). For NLP, we first demonstrate the effectiveness of C3A using the GLUE benchmark with RoBERTa-Base and RoBERTa-Large models. Additionally, we apply C3A to fine-tuning the LLaMA family models, which stands as a representative of prevalent large language models (LLMs). For CV, we examine two practical scenarios: the traditional classification task using Vision Transformers (ViTs) and the more intricate text-to-image generation task based on diffusion models.
@@ 4.1 Natural Language Understanding
Baselines. We compare our C3A with several representative PEFT methods, including BitFit Zaken et al. (2021), (IA)3 Liu et al. (2022), LoRA Hu et al. (2021), VeRA Kopiczko et al. (2023), and BOFT Liu et al. (2023). BitFit selectively fine-tunes existing parameters, specifically the biases. (IA)3 is the state-of-the-art method that adds extra adapters. LoRA is a widely known PEFT method that employs low-rank decomposition to compress additive delta matrices. VeRA is a recent approach that focuses on further reducing trainable parameters for LoRA. BOFT is another innovative method in PEFT research, compressing multiplicative delta matrices using orthogonal decomposition and butterfly factorization.
Settings. We evaluate our proposed C3A on the General Language Understanding Evaluation (GLUE) benchmark Wang et al. (2018), which encompasses a wide range of natural language understanding (NLU) tasks, including single-sentence classification, similarity and paraphrase, and natural language inference. To enhance practicality, we split these datasets following the train-validation-test approach. The best-performing model is selected based on validation set performance across the fine-tuning epochs, and the reported performance corresponds to its performance on the test set. For this evaluation, we fine-tune the pre-trained RoBERTa-Base and RoBERTa-Large models Liu et al. (2019). For the unique hyperparameters of different baselines, we adopt the values suggested in the original papers (e.g., VeRA’s $r$ and BOFT’s $b$ and $m$). The number of trainable parameters excludes the classification head, as each method uses the same one. The shared hyperparameters (i.e., the learning rate for classification head and the learning rate for other trainable parameters) are found by hyperparameter search. For the memory cost, to ensure fairness and consistency, we fix the length of input data to 256 tokens and use a batchsize of 64.
Results. Results are summarized in Table 2. In general, C3Ab=1 achieves better or on-par performance compared to baseline methods with extremely small trainable parameters. When a larger number of trainable parameters is allowed, C3Ab=6 and C3Ab=8 significantly outperform the baselines. Furthermore, compared to (IA)3, LoRA, VeRA, and BOFT, despite training a similar number of or more parameters, C3A stands out for its excellent memory efficiency. The only method with better memory efficiency is BitFit, which can be seen as an upper bound since it does not introduce any new parameters.
Table 2: Performance of different PEFT methods on the GLUE benchmark. We fine-tune pre-trained RoBERTa-Base and -Large models on 6 datasets. We report the Matthew’s correlation coefficient (MCC) for CoLA, Pearson correlation coefficient (PCC) for STS-B, and accuracy (Acc.) for all remaining tasks. For all 6 datasets, a higher metric indicates better performance. “Avg.” denotes the average accuracy of each method across all datasets. The best results for each dataset are highlighted in bold. # Trainable parameters does not include the classification head since each method uses the same head. Memory Cost is measured on fixed length (i.e., $256$) data with a batchsize of $64$.
Methods # Trainable Parameters Memory Cost (GB) SST-2 MRPC CoLA QNLI RTE STS-B Avg.
Base BitFit 102K 12.60 93.30$\pm$0.30 85.80$\pm$0.21 59.21$\pm$1.74 91.96$\pm$0.18 73.07$\pm$1.34 90.18$\pm$0.17 82.25
@@
C3Ab=1 49K 31.83 95.78$\pm$0.05 88.02$\pm$0.62 66.59$\pm$1.20 94.22$\pm$0.25 82.89$\pm$0.67 91.86$\pm$0.14 86.56
C3Ab=8 393K 31.79 95.78$\pm$0.15 88.09$\pm$0.47 67.18$\pm$1.92 94.26$\pm$0.19 84.62$\pm$1.36 91.81$\pm$0.36 86.96

### v2

@@ 3.5 Complexity Analysis
We compare the time complexity and space complexity of LoRA, VeRA and C3A in Table 1. Detailed analysis follows in this section.
@@ 3.5.1 Time Complexity
LoRA incorporates low-rank matrices $\mathbf{A}$ and $\mathbf{B}$, which are sequentially multiplied with the activation vector, resulting in a time complexity of $\mathcal{O}(r(d_{1}+d_{2}))$. Typically, $r\ll\min(d_{1},d_{2})$. Conversely, VeRA, despite its high-rank nature and minimal number of trainable parameters, incurs a prohibitive time complexity of $\mathcal{O}(r_{v}(d_{1}+d_{2}))$, where $r_{v}$ can exceed $\max(d_{1},d_{2})$. Therefore, striking a balance between high rank and computational efficiency is never a low-hanging fruit.
Thanks to the $\mathcal{O}(n\log n)$ complexity of the FFT algorithm used in Equation 1, C3A achieves a time complexity of $\mathcal{O}((d_{1}+d_{2})\log{b}+\frac{d_{1}d_{2}}{b})$. The first term is the time complexity for FFT, and the second term is for aggregation. In practical scenarios, $b$ is chosen as the greatest common divisor of $d_{1}$ and $d_{2}$ to achieve a high compression ratio. Given that, C3A is comparable to LoRA in terms of time complexity.
Table 1: Time and space complexity comparison of LoRA, VeRA and C3A. We split the space complexity into Parameter number and Other auxiliary tensors to help better understand the differences. We highlight that in practice, to achieve similar performance, $\frac{\max(d_{1},d_{2})}{b}\leq r\ll r_{v}$.
Method Time # Param # Other # Total
LoRA $\mathcal{O}(r(d_{1}+d_{2}))$ $r(d_{1}+d_{2})$ 0 $r(d_{1}+d_{2})$
VeRA $\mathcal{O}(r_{v}(d_{1}+d_{2}))$ $r_{v}+d_{1}$ $r_{v}(d_{1}+d_{2})$ $r_{v}(d_{1}+d_{2})+r_{v}+d_{1}$
C3A $\mathcal{O}((d_{1}+d_{2})\log{b}+\frac{d_{1}d_{2}}{b})$ $\frac{d_{1}d_{2}}{b}$ $b$ $\frac{d_{1}d_{2}}{b}+b$
@@ 3.5.2 Space Complexity
We analyze the space complexity of LoRA, VeRA, and C3A during training. The differences among these methods primarily arise from the trainable parameters and the auxiliary tensors required for the forward pass and backpropagation. LoRA does not rely on auxiliary tensors, while VeRA necessitates 2 random projection matrices, with a total size of $r_{v}(d_{1}+d_{2})$. Since $r_{v}$ is by no means negligible, the memory consumption of VeRA would be significantly larger than that of LoRA.
In terms of C3A, the only additional auxiliary tensor would be of size $b\leq\min(d_{1},d_{2})$, which is reserved by the FFT algorithm. Given a proper $b$, which is often close to the greatest common divisor of $d_{1}$ and $d_{2}$, the space complexity of C3A excels. Additionally, we observe that when applying FFT, input blocks and output blocks are independent within their own group. Therefore, parallelization can be employed to enable the flexibility of trading space for time.
@@ 4 Experiment
In this section, we firstly conduct experiments on a synthetic dataset to show the superior expressiveness of C3A over LoRA. After that, we comprehensively evaluate C3A in both natural language processing (NLP) and computer vision (CV). For NLP, we first demonstrate the effectiveness of C3A using the GLUE benchmark with RoBERTa-Base and RoBERTa-Large models. Additionally, we apply C3A to fine-tuning the LLaMA family models, which stands as a representative of prevalent large language models (LLMs). For CV, we examine the most representative classification task using Vision Transformers (ViTs) on a wide range of datasets.
4.1 Synthetic Data
Settings. We distribute 8 points evenly on a 2D plane as cluster centers and randomly sample 30 points from 8 corresponding multivariate Gaussian distributions. A 3-layer MLP is then used to classify these point clusters. To compare the expressiveness of 2 types of layers, we replace the middle layer with either a low-rank layer or a circulant layer, ensuring that both layers have the same number of trainable parameters for a fair comparison.
Figure 3: Expressiveness test on synthetic data. The left figure shows the synthetic data used for the experiment, while the right figure illustrates the training accuracy curves of a 3-layer MLP, incorporating C3A, LoRA, and standard linear layers, respectively.
Results. The results are presented in Figure 3. We observe that LoRAr=1 struggles with this straightforward classification task. In contrast, C3Ab=128/2, despite using the same number of parameters, achieves perfect classification, comparable to a standard linear layer. This demonstrates the high expressiveness of C3A given the same parameter budget.
@@ 4.2 Natural Language Understanding
Baselines. We compare our C3A with several representative PEFT methods, including BitFit Zaken et al. (2021), (IA)3 Liu et al. (2022), LoRA Hu et al. (2021), VeRA Kopiczko et al. (2023), and BOFT Liu et al. (2023). BitFit selectively fine-tunes existing parameters, specifically the biases. (IA)3 is the state-of-the-art method that adds extra adapters. LoRA is a widely known PEFT method that employs low-rank decomposition to compress additive delta matrices. VeRA is a recent approach that focuses on further reducing trainable parameters of LoRA while preserving a high rank. BOFT is another innovative method in PEFT research, compressing multiplicative delta matrices using orthogonal decomposition and butterfly factorization.
Table 2: Performance of different PEFT methods on the GLUE benchmark. We fine-tune pre-trained RoBERTa-Base and -Large models on 6 datasets. We report the Matthew’s Correlation Coefficient (MCC) for CoLA, Pearson Correlation Coefficient (PCC) for STS-B, and accuracy (Acc.) for all the remaining tasks. For each metric, a higher score indicates better performance. “Avg.” denotes the average score of each method across all datasets. The best results for each dataset are highlighted in bold. # Trainable parameters does not include the classification head since each method uses a head of the same size. Memory Cost is measured on fixed length (i.e., $256$) data with a batchsize of $64$.
Method # Trainable Parameters Memory Cost (GB) SST-2 MRPC CoLA QNLI RTE STS-B Avg.
Base Full 124M 17.19 94.01$\pm$0.39 87.10$\pm$0.79 62.00$\pm$1.16 92.40$\pm$0.28 77.33$\pm$2.68 90.70$\pm$0.14 83.92
BitFit 102K 12.60 93.30$\pm$0.30 85.80$\pm$0.21 59.21$\pm$1.74 91.96$\pm$0.18 73.07$\pm$1.34 90.18$\pm$0.17 82.25
@@
C3Ab=1024/1 49K 31.83 95.78$\pm$0.05 88.02$\pm$0.62 66.59$\pm$1.20 94.22$\pm$0.25 82.89$\pm$0.67 91.86$\pm$0.14 86.56
C3Ab=1024/8 393K 31.79 95.78$\pm$0.15 88.09$\pm$0.47 67.18$\pm$1.92 94.26$\pm$0.19 84.62$\pm$1.36 91.81$\pm$0.36 86.96
Settings. We evaluate our proposed C3A on the General Language Understanding Evaluation (GLUE) benchmark Wang et al. (2018), which encompasses a wide range of natural language understanding (NLU) tasks, including single-sentence classification, similarity and paraphrase, and natural language inference. To enhance practicality, we split these datasets following the train-validation-test approach. The best-performing model is selected based on validation set performance across the fine-tuning epochs, and the reported performance corresponds to its performance on the test set. For this evaluation, we fine-tune the pre-trained RoBERTa-Base and RoBERTa-Large models Liu et al. (2019). For the unique hyperparameters of different baselines, we adopt the values suggested in the original papers (e.g., VeRA’s $r$ and BOFT’s $b$ and $m$). The number of trainable parameters excludes the classification head, as each method uses one of the same size. The shared hyperparameters (i.e., the learning rate for classification head and for other trainable parameters, separately) are found by hyperparameter search. For the memory cost, to ensure fairness and consistency, we fix the length of input data to 256 tokens and use a batchsize of 64.
Results. Results are summarized in Table 2. In general, C3Ab=768/1 and C3Ab=1024/1 achieves better or on-par performance compared to baseline methods with extremely small trainable parameters. When a larger number of trainable parameters is allowed, C3Ab=768/6 and C3Ab=1024/8 significantly outperform baselines. In addition, compared to (IA)3, LoRA, VeRA, and BOFT, despite training a similar number of or more parameters, C3A stands out for its excellent memory efficiency. The only method with better memory efficiency is BitFit, which can be seen as an upper bound since it does not introduce any new parameters. Furthermore, nearly all circulant delta matrices discovered by C3A are full rank, indicating maximal capacity Zeng & Lee (2023) and providing theoretical support for the impressive results.
4.3 Instruction Tuning
Settings. We evaluate C3A against LoRA by fine-tuning LLaMA2-7B/13B and LLaMA3-8B on 4 standard instruction tuning datasets: GSM8k Cobbe et al. (2021), MATH Yu et al. (2023), ViGGO Juraska et al. (2019), and SQL Zhong et al. (2017), following the setting outlined in Nikdan et al. (2024). For the SQL dataset, we preprocess by selecting 12.5% of the dataset, followed by an 4:1 train-test split. To ensure a fair comparison, we maintain LoRA parameters with $r=32$ and a dropout rate of 0.05, while exploring various $\alpha$ values ranging from 32 to 128. All experiments are conducted for a standard single pass (epoch) over the dataset.
Results. In Table 3, we summarize our main experiments. It can be observed that C3A generally surpasses LoRA across both LLaMA2-7B/13B and LLaMA3-8B. Notably, C3A only utilizes less than half of the parameter count compared to LoRA. Overall, these results clearly highlight the effectiveness of C3A.
Table 3: Comparison of C3A and LoRA on fine-tuning LLaMA2 and LLaMA3 models in terms of accuracy and trainable parameters. The best results for each dataset are highlighted in bold. “Avg.” denotes the average accuracy of each method across all datasets.
Model Method # Trainable Parameters GSM8k MATH ViGGO SQL Avg.
LLaMA2-7B LoRAr=32 16.8M 27.22 4.10 92.24 65.39 47.24
C3Ab=4096/32 8.4M 29.34 4.25 94.03 66.21 48.46
LLaMA2-13B LoRAr=32 26.2M 42.68 6.70 93.19 67.53 52.53
C3Ab=5120/32 13.1M 42.00 6.65 95.57 67.28 52.88
LLaMA3-8B LoRAr=32 13.6M 56.57 18.95 91.78 68.19 58.87
C3Ab=4096/32 5.2M 57.54 19.75 93.07 69.47 59.96
4.4 Image Classification
Settings. We focus on the task of image classification utilizing the Vision Transformer (ViT) models. Specifically, we employ both the Base and Large variants of this prevalent computer vision foundation model, as introduced by Dosovitskiy et al. (2020). These ViT models are pre-trained on the extensive ImageNet-21K dataset Ridnik et al. (2021). For the fine-tuning phase, we utilize a diverse set of datasets including Pets, Cars, DTD, EuroSAT, FGVC and RESISC.
Table 4: Fine-tuning results with ViT-Base and ViT-Large models on various image classification datasets. The models are fine-tuned for 10 epochs, and the best-performing model, based on validation set accuracy, is selected. The reported accuracy corresponds to the performance on the test set. The best results for each dataset are highlighted in bold. “Avg.” denotes the average accuracy of each method across all datasets.
Method # Trainable Parameters Pets Cars DTD EuroSAT FGVC RESISC Avg.
Base Head - 90.28$\pm$0.43 25.76$\pm$0.28 69.77$\pm$0.67 88.72$\pm$0.13 17.44$\pm$0.43 74.22$\pm$0.10 61.03
Full 85.8M 92.82$\pm$0.54 85.10$\pm$0.21 80.11$\pm$0.56 99.11$\pm$0.07 61.60$\pm$1.00 96.00$\pm$0.23 85.79
LoRAr=16 590K 93.76$\pm$0.44 78.04$\pm$0.33 78.56$\pm$0.62 98.84$\pm$0.08 56.64$\pm$0.55 94.66$\pm$0.17 83.42
C3Ab=768/12 221K 93.88$\pm$0.22 79.05$\pm$0.35 80.57$\pm$0.53 98.88$\pm$0.07 54.31$\pm$0.79 94.54$\pm$0.23 83.54
Large Head - 91.11$\pm$0.30 37.91$\pm$0.27 73.33$\pm$0.26 92.64$\pm$0.08 24.62$\pm$0.24 82.02$\pm$0.11 66.94
Full 303M 94.30$\pm$0.31 88.15$\pm$0.50 80.18$\pm$0.66 99.06$\pm$0.10 67.38$\pm$1.06 96.08$\pm$0.20 87.53
LoRAr=16 1.57M 94.62$\pm$0.47 86.11$\pm$0.42 80.09$\pm$0.42 98.99$\pm$0.03 63.64$\pm$0.83 95.94$\pm$0.21 86.56
C3Ab=1024/16 786K 94.48$\pm$0.30 84.94$\pm$0.39 82.62$\pm$0.52 98.75$\pm$0.19 63.80$\pm$0.37 95.52$\pm$0.16 86.69
Results. Table 4 presents a summary of the results obtained from 6 image classification datasets using the ViT Base and Large models. The LoRA and C3A methods both show substantial performance improvements over Head Tuning, highlighting their effectiveness in the image classification domain. Notably, our method achieves comparable performance to LoRA while utilizing only half of the parameter count.